10. A car is being driven along a straight, level highway at 71.582 mph. The driver taps on the brakes but notices no response. He tries to brake again and realizes the brakes are completely out. There is a field next to the highway with tall grass. The driver veers into the tall grass traveling in a straight line. The car eventually comes to a stop. Later it is measured that the car traveled 184 meters through the grass. How much time did it take for the vehicle to come to a stop after leaving the road? Answer: 11.5 seconds
Answer #1

initial velocity = V0 = 71. 582mph = 32 m/sec

displacement = S = 184 m

final velocity = Vf = 0 m/sec ( stop)

time = t = ?

displacement = Avearage velocity x time

S = ( VF + V0 ) / 2 * T

184 = (32 / 2 ) * t

time = t = 184 x 2 / 32 = 11.5 sec